The Minot Kit Aircraft Company of​ Minot, North​ Dakota, uses a plasma cutter to fabricate metal aircraft parts for its plane kits. The company currently is using a cutter that it purchased four years ago that has a book value of ​$60,000 and is being depreciated ​$15,000 per year over the next 4 years. If the old cutter were to be sold​ today, the company estimates that it would bring in an amount equal to the book value of the equipment. The company is considering the purchase of a new automated plasma cutter that would cost ​$300,000 to install and that would be depreciated over the next 4 years toward a ​$44,000 salvage value using​ straight-line depreciation. The primary advantage of the new cutter is the fact that it is fully automated and can be run by one operator rather than the three employees that are currently required. The labor savings would be ​$80,000 per year. The firm faces an income tax rate of 28 percent. At the end of the​ 4-year project both cutters could be sold at their​ end-of-project book value.

a.  What are the differential operating cash flow savings per year during years 1 through 4 for the new plasma​ cutter?

b.  What is the initial cash outlay required to replace the existing plasma cutter with the newer​ model?

c.  What does the timeline for the replacement project cash flows for years 0 through 4 look​ like?

d.  If the company requires a discount rate of 9 percent for new​ investments, should the plasma cutter be​ replaced?
Answer #1

a). Depreciation of old machine (D1) per year = 15,000

Depreciation of new machine (D2) per year = (300,000 - 44,000)/4 = 64,000

Incremental depreciation (ID) per year = D2 - D1 = 64,000 - 15,000 = 49,000

Differential operating cash flow (OCF) = Labor savings*(1-Tax rate) + (ID*Tax rate) = 80,000*(1-28%) + (49,000*28%) = 71,320

b). Initial cash outlay = purchase price of new machine - opportunity cost of old machine

= 300,000 - 60,000 = 240,000

c & d). Replacement project cash flows and project NPV are calculated in the table below:

-240000 SV Formula Time (n) co Initial cash outlay OCF Incremental OCF Salvage value CF = OCF + SV - CO Project cash flow 1/(

As project NPV is positive, the cutter can be replaced.